OpenStudy (anonymous):

Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and distance traveled by the toy car. (4 points) [Choose the value of correlation coefficient from 1, 0.99, 0.5, 0.02]

OpenStudy (anonymous):

Time (x) (minutes) 10 20 30 40 Distance (y) (miles) 4 8 12 15

OpenStudy (cj49):

4/10 = 0.4 miles/min 8/20 = 0.4 12/30=0.4 15/40=0.4 if they are then the correlation coefficient is 1